Calcute the pH of a 0.0001 M KOH solution?
1
Expert's answer
2013-03-07T10:11:39-0500
pH = -lg [H+]

In a water solution the ion product for water is:
[H+][OH-] = Kw= 1 x 10-14

lg [H+] + lg [OH-] = lg[10-14]

pH + pOH = 14

pH – lg[OH-] = 14

pH = 14 + lg[OH-]

c(OH-) = c(KOH) = 0.0001 M

pH = 14 + lg 0.0001 = 10

Answer: pH = 10